OpenStudy (anonymous):

@jim_thompson5910 give me a sec please

OpenStudy (anonymous):

ABCD is a square with a side length of 9 units. If the ratio of CD:CE is 2:5, what is the area of ABCDE? I'll draw it below

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

how long is DE or ED

OpenStudy (anonymous):

I'm not given that

jimthompson5910 (jim_thompson5910):

you can easily figure it out though hint: ABCD is a square

jimthompson5910 (jim_thompson5910):

sorry I meant to say "ABDE is a square"

OpenStudy (anonymous):

oh it'll be 9!

jimthompson5910 (jim_thompson5910):

yes, so if CD = x, then we can set up this ratio CD/CE = 2/5 CD/(DE+CD) = 2/5 x/(9+x) = 2/5 5x = 2(9+x) now solve for x

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now you need to find the area of the square and the area of the triangle

OpenStudy (anonymous):

The area of the square will be 6 and I think the area of the triangle will be 15 ?

jimthompson5910 (jim_thompson5910):

the area of the square is not 6

jimthompson5910 (jim_thompson5910):

remember, the sides of the square are 9

OpenStudy (anonymous):

oh no! i mean 81, my bad!

jimthompson5910 (jim_thompson5910):

good, the area of the triangle is not 15

jimthompson5910 (jim_thompson5910):

the base is 6, the height is 9 use that in A = b*h/2

OpenStudy (anonymous):

54

jimthompson5910 (jim_thompson5910):

now cut that in half

OpenStudy (anonymous):

27

jimthompson5910 (jim_thompson5910):

good, so the total area is: 81+27 = 108

jimthompson5910 (jim_thompson5910):

the units are "square units"

OpenStudy (anonymous):

I'm confused on how x=6 relates back to the problem

jimthompson5910 (jim_thompson5910):

x = 6 is the base of the triangle

jimthompson5910 (jim_thompson5910):

it's the length of CD

OpenStudy (anonymous):

oh i taught the base was 2 bc on the problem it says CD:C3 is 2:5

jimthompson5910 (jim_thompson5910):

that 2 is the ratio, it's not the actual length of CD

jimthompson5910 (jim_thompson5910):

part of the ratio I should say

OpenStudy (anonymous):

I think I have a better understanding of it now! Thank you oso much! I really mean it

jimthompson5910 (jim_thompson5910):

you're welcome